# SAMSGL: Series-Aligned Multi-Scale Graph Learning for Spatio-Temporal Forecasting

### v2

@@ III-B Series-Aligned Graph Convolution
where $\mathscr{F}(\bar{Q})$ entails a fast Fourier transform applied to $\bar{Q}$. *$\mathscr{F}^{*}(K)$ corresponds to the conjugation of the fast Fourier transform of $K$, and $\mathscr{F}^{-1}(\cdot)$ refers to the inversed fast Fourier transform. After deriving correlation scores at each time step between nodes in the graph and the reference node, we identify the time step with the highest correlation score as the time of messages arrival, since nodes often exhibit similar patterns within message propagation through a graph. Consequently, we determine the time delay between the reference node and each node $i$ as follows:
@@
where $\tilde{X}_{i,t}$ corresponds to the time series after alignment of node $i$ at time step $t$. Then the aligned time series $\tilde{X}$ are reweighed with correlation scores $\zeta$, and a $1$-dimensional convolution layer denoted as “Conv1d” is utilized. Subsequently, we proceed to execute the graph convolution operation on the adjacent matrix $\mathcal{A}_{\rm nd}$ of the non-delayed graph.
@@
III-C Multi-Scale Graph Structure Learning
The high-dimensional interactions between nodes can be characterized by connections among a large number of nodes across various spatial scales. To learn spatial interactions, we adopt multi-scale graph structure learning for different spatial scales. We construct graphs in a multi-scale sense, including the global delayed and non-delayed graphs, as well as the local graph. These two global graphs depict global interactions from two pertinent perspectives based on the overall node topology, which are tailored for spatial graph aggregation and aligned node features. Inspired by [16] which employs a shared network to ensure the universality of graph structures, we generate the global delayed and non-delayed graphs employing separate trainable node embeddings and a shared network $\mathcal{F}_{E}$. These embeddings are generated randomly to ensure diverse graph perspectives. The procedure unfolds as follows:
@@ III-C Multi-Scale Graph Structure Learning
where $c$ denotes the delayed or non-delayed classes, and $c\in{\{\rm d,\rm nd\}}$. $E_{c}$ is the trainable node embedding. We derive the generalized node embedding $\tilde{E}_{c}$ by feeding $E_{c}$ into the shared network $\mathcal{F}_{E}$, and $\tilde{E}_{c}^{T}$ is the transpose of $\tilde{E}_{c}$. $\hat{\mathcal{A}}_{c}$ is the raw adjacent matrix formed using node embeddings. To prevent over-fitting and maintain graph sparsity, we implement a regularization operation [30] on the raw adjacent matrix.
@@
where $\sigma$ is the activation function, and $\hat{a}_{c,i,j}$ is an element of $\hat{\mathcal{A}}_{c}$. $g_{i,j}$ and $g_{i,j}^{{}^{\prime}}$ are drawn from a Gumbel(0,1) distribution, of which the probability density function is $f(z;\mu,\beta)=\frac{1}{\beta}\exp\left(-\frac{z-\mu}{\beta}-\exp(-(z-\mu)/\beta)\right)$. $z$ represents the random variable, $\mu=0$ is the location parameter and $\beta=1$ is the scale parameter. $s$ is a hyper-parameter related to temperature. $a_{c,i,j}$ is an element of the regularized adjacent matrix $\mathcal{A}_{c}$. By implementing regularization to $\hat{\mathcal{A}}_{\rm nd}$ and $\hat{\mathcal{A}}_{\rm d}$, we derive the adjacent matrices $\mathcal{A}_{\rm nd}$ and $\mathcal{A}_{\rm d}$ of global non-delayed and delayed graphs.
@@
where ${a}_{{\rm l},i,j}$ represents an element of the adjacent matrix $\mathcal{A}_{\rm l}$ of local graph. $d_{i,j}$ corresponds to the distance between node $i$ and node $j$. $\alpha$ is a hyper-parameter to scale the distance, and $r_{i,j}$ signifies the relationship between node $i$ and node $j$, which may encompass relative coordinates like $y_{i}$ and $y_{j}$ or distance $d_{i,j}$ between node $i$ and node $j$.
It is worth noting that we remap locations distributed across a spherical surface and incorporate the angles between nodes [8] to calculate the adjacent matrix for large-scale global spatio-temporal forecasting tasks. Through the process of global and local graph structure learning, we acquire $\mathcal{A}_{\rm nd}$ of the global non-delayed graph, $\mathcal{A}_{\rm d}$ of the global delayed graph and $\mathcal{A}_{\rm l}$ of the local graph.
III-D Graph-FC Blocks
To capture spatio-temporal dependencies and fuse node features from multi-scale graphs, we build Graph-FC blocks. These blocks combine multi-graph convolution layers with backward-forward residual structures to iteratively learn spatio-temporal dynamics. As shown in FI 1 (b), each Graph-FC block utilizes multi-graph convolution to capture features from multi-scale graphs. The outputs of graph convolution, $H_{\rm nd}$, $H_{\rm d}$, and $H_{\rm l}$, are then concatenated and fed into a fully-connected layer. The multi-graph convolution, which captures node features from global delayed, non-delayed, and local scales, can be expressed as follows:
@@ III-D Graph-FC Blocks
where $X^{m}$ represents the input for the $m$-th block. $W_{\rm d}$ and $W_{\rm l}$ are the trainable parameters corresponding to delayed graph convolution and local graph convolution. The subsequent step involves reshaping the hidden state of the $k$-th block $H^{m}_{0}\in\mathbb{R}^{{L}\times{N}\times{D}}$ into $H^{m}_{1}\in\mathbb{R}^{{N}\times{(L\times D)}}$ and applying the backward-forward residual operation. Following this, as shown in FI 1 (a), we pass the hidden state through a MLP to capture temporal features $H_{2}^{m}$. Finally, two fully-connected layers ${\rm FC}_{b}$ and ${\rm FC}_{f}$ are employed to map $H^{m}_{2}$ to historical and forecasting lengths.
@@ IV Experiments
TABLE I: Baseline methods.
Classes Baseline Methods References
RNN-based models Graph Convolutional Gate Recurrent Unit (GCGRU) [43]
Diffusion Convolutional Recurrent Neural Network (DCRNN) [24]
Temporal Graph Convolutional Network (TGCN) [25]
Adaptive Graph Convolutional Recurrent Network (AGCRN) [13]
Conditional Local Convolution Recurrent Network (CLCRN) [8]
Scalable Graph Predictor (SGP) [50]
Spatial and temporal attention models Spatio-Temporal Graph Convolutional Network (STGCN) [44]
Multi-Component Spatial-Temporal Graph Convolution Networks (MSTGCN) [18]
Attention based Spatio-Temporal Graph Convolutional Networks (ASTGCN) [18]
Graph structure learning models Meta-Graph Convolutional Recurrent Network (MGCRN) [16]
Regularized Graph Structure Learning (RGSL) [30]
Transformer-based model Corrformer [9]
@@
TABLE II: Overall forecasting performance of different methods on different meteorology and traffic datasets.
@@
    Methods & Metrics Datasets Wind Temperature Cloud cover Humidity PeMSD4 PeMSD8    TGCN [25] MAE 4.1747$\pm$0.0324 3.8638$\pm$0.0970 2.3934$\pm$0.0216 1.4700$\pm$0.0295 34.7859$\pm$0.2043 38.2618$\pm$1.0967 RMSE 5.6730$\pm$0.0412 5.8554$\pm$0.1432 3.6512$\pm$0.0223 2.1066$\pm$0.0551 52.4670$\pm$0.2708 57.0713$\pm$1.5649    STGCN [44] MAE 3.6477$\pm$0.0000 4.3525$\pm$1.0442 2.019$\pm$0.0392 0.7975$\pm$0.2378 25.3017$\pm$2.3161 21.3431$\pm$0.1006 RMSE 4.8146$\pm$0.0003 6.8600$\pm$1.1233 2.9542$\pm$0.0542 1.1109$\pm$0.2913 38.9132$\pm$2.2506 32.5995$\pm$0.2007    MSTGCN [18] MAE 1.9440$\pm$0.0150 1.2199$\pm$0.0058 1.8732$\pm$0.0010 0.6093$\pm$0.0012 21.9152$\pm$0.6523 18.5269$\pm$0.2870 RMSE 2.9111$\pm$0.0292 1.9203$\pm$0.0093 2.8629$\pm$0.0073 0.8684$\pm$0.0019 34.6353$\pm$0.9230 28.5750$\pm$0.4316    ASTGCN [18] MAE 2.0889$\pm$0.0006 1.4896$\pm$0.0130 1.9936$\pm$0.0002 0.7288$\pm$0.0229 23.5648$\pm$0.9421 18.4513$\pm$0.5240 RMSE 3.1356$\pm$0.0012 2.4622$\pm$0.0023 2.9576$\pm$0.0007 1.0471$\pm$0.0402 36.7932$\pm$1.3277 28.4153$\pm$0.8141    GCGRU [43] MAE 1.4116$\pm$0.0057 1.3256$\pm$0.1499 1.5938$\pm$0.0021 0.5007$\pm$0.0002 25.8336$\pm$0.0399 20.2753$\pm$0.0572 RMSE 2.2931$\pm$0.0047 2.1721$\pm$0.1945 2.5576$\pm$0.0116 0.7891$\pm$0.0006 39.8696$\pm$0.1136 31.8069$\pm$0.0812    DCRNN [24] MAE 1.4321$\pm$0.0019 1.3232$\pm$0.0864 1.5938$\pm$0.0021 0.5046$\pm$0.0011 24.9117$\pm$2.0638 20.2303$\pm$0.1051 RMSE 2.3364$\pm$0.0055 2.1874$\pm$0.1227 2.5412$\pm$0.0044 0.7956$\pm$0.0033 38.619$\pm$2.8952 31.7793$\pm$0.0376    AGCRN [13] MAE 2.4194$\pm$0.1149 1.2651$\pm$0.0080 1.7501$\pm$0.1467 0.5759$\pm$0.1632 19.3291$\pm$0.3053 15.8135$\pm$0.0957 RMSE 3.4171$\pm$0.0055 1.9314$\pm$0.0219 2.7585$\pm$0.1694 0.8549$\pm$0.2025 31.3227$\pm$0.5927 25.5016$\pm$0.1877    CLCRN [8] MAE 1.3260$\pm$0.0483 1.1688$\pm$0.0457 1.4906$\pm$0.0037 0.4531$\pm$0.0065 127.5582$\pm$0.1709 120.7720$\pm$0.0037 RMSE 2.129$\pm$0.0733 1.8825$\pm$0.1509 2.4559$\pm$0.0037 0.7078$\pm$0.0146 159.9446$\pm$0.7968 146.6032$\pm$0.0159    Corrformer [9] MAE 1.3157$\pm$0.0007 0.8055$\pm$0.0014 0.1559$\pm$0.0001 4.6720$\pm$0.0023 25.4462$\pm$0.1916 20.2820$\pm$0.1950 RMSE 2.1507$\pm$0.0007 1.3627$\pm$0.0011 0.2473$\pm$0.0005 7.1811$\pm$0.0041 37.7988$\pm$0.2390 30.1798$\pm$0.2621    MGCRN [16] MAE 1.3899$\pm$0.0035 1.1143$\pm$0.0709 1.5930$\pm$0.0035 0.4481$\pm$0.0012 18.8580$\pm$0.0413 15.2828$\pm$0.0410 RMSE 2.2696$\pm$0.0099 1.8806$\pm$0.1187 2.5409$\pm$0.0154 0.6903$\pm$0.0043 30.9136$\pm$0.1235 24.6436$\pm$0.2891    RGSL [30] MAE 1.1927$\pm$0.0096 0.7877$\pm$0.0353 1.4911$\pm$0.0015 0.4165$\pm$0.0072 19.5440$\pm$0.2571 15.7333$\pm$0.2037 RMSE 1.9342$\pm$0.0161 1.3233$\pm$0.0397 2.4136$\pm$0.0052 0.6385$\pm$0.0097 31.0884$\pm$0.1087 25.0912$\pm$0.2831    Ours MAE 0.9014$\pm$0.0091 0.6329$\pm$0.0131 1.4086$\pm$0.0037 0.3761$\pm$0.0018 19.1346$\pm$0.0627 15.2174$\pm$0.0035 RMSE 1.4072$\pm$0.0157 1.0876$\pm$0.0194 2.3084$\pm$0.0084 0.5807$\pm$0.0028 30.6541$\pm$0.0776 24.4680$\pm$0.0015

### v3

@@ III.2 Series-Aligned Graph Convolution
where $\mathscr{F}(\bar{Q})$ entails a fast Fourier transform applied to $\bar{Q}$. $\mathscr{F}^{*}(K)$ corresponds to the conjugation of the fast Fourier transform of $K$, and $\mathscr{F}^{-1}(\cdot)$ refers to the inversed fast Fourier transform. After deriving correlation scores at each time step between nodes in the graph and the reference node, we identify the time step with the highest correlation score as the time of messages arrival, since nodes often exhibit similar patterns within message propagation through a graph. Consequently, we determine the time delay between the reference node and each node $i$ as follows:
@@
where $\tilde{X}_{i,t}$ corresponds to the time series after alignment of node $i$ at time step $t$. Then the aligned time series $\tilde{X}$ is reweighed with correlation scores $\zeta$, and a $1$-dimensional convolution layer denoted as “Conv1d" is utilized. Subsequently, we proceed to execute the graph convolution operation on the adjacency matrix $\mathcal{A}_{\rm nd}$ of the non-delayed graph.
@@
III.3 Multi-Scale Graph Structure Learning
The high-dimensional interactions between nodes can be characterized by connections among a large number of nodes across various spatial scales. To learn spatial interactions, we adopt multi-scale graph structure learning for different spatial scales. We construct graphs in a multi-scale sense, including the global delayed and non-delayed graphs, as well as the local graph. These two global graphs depict global interactions from two pertinent perspectives based on the overall node topology, which are tailored for spatial graph aggregation and aligned node features. Inspired by Jiang et al.Jiang et al. (2023a) which employ a shared network to ensure the universality of graph structures, we generate the global delayed and non-delayed graphs employing separate trainable node embeddings and a shared network $\mathcal{F}_{E}$. These embeddings are generated randomly to ensure diverse graph perspectives. The procedure unfolds as follows:
@@ III.3 Multi-Scale Graph Structure Learning
where $c$ denotes the delayed or non-delayed classes, and $c\in{\{\rm d,\rm nd\}}$. $E_{c}$ is the trainable node embedding. We derive the generalized node embedding $\tilde{E}_{c}$ by feeding $E_{c}$ into the shared network $\mathcal{F}_{E}$, and $\tilde{E}_{c}^{T}$ is the transpose of $\tilde{E}_{c}$. $\hat{\mathcal{A}}_{c}$ is the raw adjacency matrix formed using node embeddings. To prevent over-fitting and maintain graph sparsity, we implement a regularization operation Yu et al. (2022) on the raw adjacency matrix.
@@
where $\sigma$ is the activation function, and $\hat{a}_{c,i,j}$ is an element of $\hat{\mathcal{A}}_{c}$. $g_{i,j}$ and $g_{i,j}^{{}^{\prime}}$ are drawn from a Gumbel(0,1) distribution, of which the probability density function is $f(z;\mu,\beta)=\frac{1}{\beta}\exp\left(-\frac{z-\mu}{\beta}-\exp(-(z-\mu)/\beta)\right)$. $z$ represents the random variable, $\mu=0$ is the location parameter and $\beta=1$ is the scale parameter. $s$ is a hyper-parameter related to temperature. $a_{c,i,j}$ is an element of the regularized adjacency matrix $\mathcal{A}_{c}$. By implementing regularization to $\hat{\mathcal{A}}_{\rm nd}$ and $\hat{\mathcal{A}}_{\rm d}$, we derive the adjacency matrices $\mathcal{A}_{\rm nd}$ and $\mathcal{A}_{\rm d}$ of global non-delayed and delayed graphs.
@@
where ${a}_{{\rm l},i,j}$ represents an element of the adjacency matrix $\mathcal{A}_{\rm l}$ of local graph. $d_{i,j}$ corresponds to the distance between node $i$ and node $j$. $\alpha$ is a hyper-parameter to scale the distance. The value $r_{i,j}$ is derived from the relative location computing module, which captures the relationship between node $i$ and node $j$. This relationship may encompass relative coordinates such as $y_{i}$ and $y_{j}$, or the distance $d_{i,j}$ between the nodes, depending on the available spatial information.
It is worth noting that we remap locations distributed across a spherical surface and incorporate the angles between nodes Lin et al. (2022) to calculate the adjacency matrix for large-scale global spatio-temporal forecasting tasks. Through the process of global and local graph structure learning, we acquire $\mathcal{A}_{\rm nd}$ of the global non-delayed graph, $\mathcal{A}_{\rm d}$ of the global delayed graph and $\mathcal{A}_{\rm l}$ of the local graph.
III.4 Graph-FC Blocks
To capture spatio-temporal dependencies and fuse node features from multi-scale graphs, we build Graph-FC blocks. These blocks combine multi-graph convolution layers with backward-forward residual structures to iteratively learn spatio-temporal dynamics. As shown in Fig. 1 (b), each Graph-FC block utilizes multi-graph convolution to capture features from multi-scale graphs. The outputs of graph convolution, $H_{\rm nd}$, $H_{\rm d}$, and $H_{\rm l}$, are then concatenated and fed into a fully-connected layer. The multi-graph convolution, which captures node features from global delayed, non-delayed, and local scales, can be expressed as follows:
@@ III.4 Graph-FC Blocks
Table 1: Baseline methods.
Classes Baseline Methods Graph structure learning
RNN-based models Graph Convolutional Gate Recurrent Unit (GCGRU) Seo et al. (2018)
Diffusion Convolutional Recurrent Neural Network (DCRNN) Li et al. (2018)
Temporal Graph Convolutional Network (TGCN) Zhao et al. (2020)
Adaptive Graph Convolutional Recurrent Network (AGCRN) Bai et al. (2020)
Conditional Local Convolution Recurrent Network (CLCRN) Lin et al. (2022)
Scalable Graph Predictor (SGP) Cini et al. (2023)
Meta-Graph Convolutional Recurrent Network (MGCRN) Jiang et al. (2023a)
Regularized Graph Structure Learning (RGSL) Yu et al. (2022)
Time-space alternation Spatio-Temporal Graph Convolutional Network (STGCN) Yu, Yin, and Zhu (2018)
Multi-Component Spatial-Temporal Graph Convolution Networks (MSTGCN) Guo et al. (2019)
Attention based Spatio-Temporal Graph Convolutional Networks (ASTGCN) Guo et al. (2019)
Transformer-based model Corrformer Wu et al. (2023)
where $X^{m}$ represents the input for the $m$-th block. $W_{\rm d}$ and $W_{\rm l}$ are the trainable parameters corresponding to delayed graph convolution and local graph convolution. The subsequent step involves reshaping the hidden state of the $k$-th block $H^{m}_{0}\in\mathbb{R}^{{L}\times{N}\times{D}}$ into $H^{m}_{1}\in\mathbb{R}^{{N}\times{(L\times D)}}$ and applying the backward-forward residual operation. Following this, as shown in Fig. 1 (a), we pass the hidden state through a MLP to capture temporal features $H_{2}^{m}$. Finally, two fully-connected layers ${\rm FC}_{b}$ and ${\rm FC}_{f}$ are employed to map $H^{m}_{2}$ to historical and forecasting lengths.
@@ IV Experiments
Table 2: Overall forecasting performance of different methods on different meteorology and traffic datasets.
@@
Methods & Metrics Datasets Wind Temperature Cloud cover Humidity PeMSD4 PeMSD8
TGCN Zhao et al. (2020) MAE 4.1747$\pm$0.0324 3.8638$\pm$0.0970 2.3934$\pm$0.0216 1.4700$\pm$0.0295 34.7859$\pm$0.2043 38.2618$\pm$1.0967
RMSE 5.6730$\pm$0.0412 5.8554$\pm$0.1432 3.6512$\pm$0.0223 2.1066$\pm$0.0551 52.4670$\pm$0.2708 57.0713$\pm$1.5649
STGCN Yu, Yin, and Zhu (2018) MAE 3.6477$\pm$0.0000 4.3525$\pm$1.0442 2.019$\pm$0.0392 0.7975$\pm$0.2378 25.3017$\pm$2.3161 21.3431$\pm$0.1006
RMSE 4.8146$\pm$0.0003 6.8600$\pm$1.1233 2.9542$\pm$0.0542 1.1109$\pm$0.2913 38.9132$\pm$2.2506 32.5995$\pm$0.2007
MSTGCN Guo et al. (2019) MAE 1.9440$\pm$0.0150 1.2199$\pm$0.0058 1.8732$\pm$0.0010 0.6093$\pm$0.0012 21.9152$\pm$0.6523 18.5269$\pm$0.2870
RMSE 2.9111$\pm$0.0292 1.9203$\pm$0.0093 2.8629$\pm$0.0073 0.8684$\pm$0.0019 34.6353$\pm$0.9230 28.5750$\pm$0.4316
ASTGCN Guo et al. (2019) MAE 2.0889$\pm$0.0006 1.4896$\pm$0.0130 1.9936$\pm$0.0002 0.7288$\pm$0.0229 23.5648$\pm$0.9421 18.4513$\pm$0.5240
RMSE 3.1356$\pm$0.0012 2.4622$\pm$0.0023 2.9576$\pm$0.0007 1.0471$\pm$0.0402 36.7932$\pm$1.3277 28.4153$\pm$0.8141
GCGRU Seo et al. (2018) MAE 1.4116$\pm$0.0057 1.3256$\pm$0.1499 1.5938$\pm$0.0021 0.5007$\pm$0.0002 25.8336$\pm$0.0399 20.2753$\pm$0.0572
RMSE 2.2931$\pm$0.0047 2.1721$\pm$0.1945 2.5576$\pm$0.0116 0.7891$\pm$0.0006 39.8696$\pm$0.1136 31.8069$\pm$0.0812
DCRNN Li et al. (2018) MAE 1.4321$\pm$0.0019 1.3232$\pm$0.0864 1.5938$\pm$0.0021 0.5046$\pm$0.0011 24.9117$\pm$2.0638 20.2303$\pm$0.1051
RMSE 2.3364$\pm$0.0055 2.1874$\pm$0.1227 2.5412$\pm$0.0044 0.7956$\pm$0.0033 38.619$\pm$2.8952 31.7793$\pm$0.0376
AGCRN Bai et al. (2020) MAE 2.4194$\pm$0.1149 1.2651$\pm$0.0080 1.7501$\pm$0.1467 0.5759$\pm$0.1632 19.3291$\pm$0.3053 15.8135$\pm$0.0957
RMSE 3.4171$\pm$0.0055 1.9314$\pm$0.0219 2.7585$\pm$0.1694 0.8549$\pm$0.2025 31.3227$\pm$0.5927 25.5016$\pm$0.1877
CLCRN Lin et al. (2022) MAE 1.3260$\pm$0.0483 1.1688$\pm$0.0457 1.4906$\pm$0.0037 0.4531$\pm$0.0065 127.5582$\pm$0.1709 120.7720$\pm$0.0037
RMSE 2.129$\pm$0.0733 1.8825$\pm$0.1509 2.4559$\pm$0.0037 0.7078$\pm$0.0146 159.9446$\pm$0.7968 146.6032$\pm$0.0159
Corrformer Wu et al. (2023) MAE 1.3157$\pm$0.0007 0.8055$\pm$0.0014 0.1559$\pm$0.0001 4.6720$\pm$0.0023 25.4462$\pm$0.1916 20.2820$\pm$0.1950
RMSE 2.1507$\pm$0.0007 1.3627$\pm$0.0011 0.2473$\pm$0.0005 7.1811$\pm$0.0041 37.7988$\pm$0.2390 30.1798$\pm$0.2621
MGCRN Jiang et al. (2023a) MAE 1.3899$\pm$0.0035 1.1143$\pm$0.0709 1.5930$\pm$0.0035 0.4481$\pm$0.0012 18.8580$\pm$0.0413 15.2828$\pm$0.0410
RMSE 2.2696$\pm$0.0099 1.8806$\pm$0.1187 2.5409$\pm$0.0154 0.6903$\pm$0.0043 30.9136$\pm$0.1235 24.6436$\pm$0.2891
RGSL Yu et al. (2022) MAE 1.1927$\pm$0.0096 0.7877$\pm$0.0353 1.4911$\pm$0.0015 0.4165$\pm$0.0072 19.5440$\pm$0.2571 15.7333$\pm$0.2037
RMSE 1.9342$\pm$0.0161 1.3233$\pm$0.0397 2.4136$\pm$0.0052 0.6385$\pm$0.0097 31.0884$\pm$0.1087 25.0912$\pm$0.2831
Ours MAE 0.9014$\pm$0.0091 0.6329$\pm$0.0131 1.4086$\pm$0.0037 0.3761$\pm$0.0018 19.1346$\pm$0.0627 15.2174$\pm$0.0035
RMSE 1.4072$\pm$0.0157 1.0876$\pm$0.0194 2.3084$\pm$0.0084 0.5807$\pm$0.0028 30.6541$\pm$0.0776 24.4680$\pm$0.0015